# WF-VAE: Enhancing Video VAE by Wavelet-Driven Energy Flow for Latent Video Diffusion Model

### v2

@@ 1 Introduction
The release of Sora [5], a video generation model developed by OpenAI, has pushed the boundaries of synthesizing photorealistic videos, drawing unprecedented attention to the field of video generation. Recent advancements in Latent Video Diffusion Models (LVDMs), such as Open-Sora Plan [19], Open-Sora [45], CogVideoX [39], EasyAnimate [38], Movie Gen [25], and Allegro [46], have led to substantial improvements in video generation quality. These methods establish a compressed latent space [27] using a pretrained video Variational Autoencoder (VAE), where the generative performance is fundamentally determined by the compression quality.
@@
Current video VAEs remain constrained by fully convolutional architectures inherited from image era. They attempt to address video flickering and redundant information by incorporating spatio-temporal interaction layers and spatio-temporal compression layers. Several recent works, including OD-VAE [19, 6], CogVideoX [39], CV-VAE [44], and Allegro [46] adopt dense 3D structure to achieve high-quality video compression. While these methods demonstrate impressive reconstruction performance, they require prohibitively intensive computational resources. In contrast, alternative approaches such as Movie Gen [25], and Open-Sora [45] utilize 2+1D architecture, resulting in reduced computational requirements at the cost of lower reconstruction quality. Previous architectures have not fully exploited video temporal redundancy, leading to requirements for redundant spatio-temporal interaction layers to enhance video compression quality.
@@
Many attempts employ block-wise inference strategies to trade computation time for memory, thus addressing computational bottlenecks in processing high-resolution, long-duration videos. EasyAnimate [38] introduced Slice VAE to encode and decode video frames in groups, but this approach leads to discontinuous output videos. Several methods, including Open-Sora [45], Open-Sora Plan [19], Allegro [46], and Movie Gen [25], implement tiling inference strategies, but often produce spatio-temporal artifacts in overlapping regions. Although CogVideoX [39] employ caching to ensure convolution continuity, its reliance on group normalization [36] disrupts the independence of temporal feature, thus preventing lossless block-wise inference. These limitations highlight two core challenges faced by video VAEs: (1) excessive computational demands due to redundant architecture, and (2) compromised latent space integrity resulting from existing tiling inference strategies, which cause artifacts and flickering in reconstructed videos.
@@
Wavelet transform [23, 26] decomposes videos into multiple frequency-domain components. This decomposition enables prioritization strategies for encoding crucial video components. In this work, we propose Wavelet Flow VAE (WF-VAE), a novel autoencoder that utilizes multi-level wavelet transforms for extracting multi-scale pyramidal features and establishes a main energy flow pathway for these features to flow into latent representation. This pathway bypasses low-frequency video information to latent space, skipping the backbone network. Our WF-VAE enables a simplified backbone design with reduced 3D convolutions, significantly reducing computational costs. To address the potential latent space disruption, we propose Causal Cache mechanism. This approach leverages the properties of causal convolution and maintains the continuity of the convolution sliding window through a caching strategy, which ensuring numerical identity between block-wise inference and direct inference results. Experimental results show that WF-VAE achieves state-of-the-art performance in both reconstruction quality and computational efficiency. To summarize, major contributions of our work include:
@@
  • We introduce a lossless block-wise inference mechanism called Causal Cache which maintains identical performance as direct inference across videos of any duration.
  • Extensive experimental evaluations of video reconstruction and generation demonstrate that WF-VAE achieves state-of-the-art performance, as validated through comprehensive ablation studies.
@@ 2 Related Work
Figure 2: Overview of WF-VAE. Our architecture consists of a backbone and a main energy flow pathway. The pathway functions as a “highway” for main flow of video energy, channeling this energy into the backbone through concatenations, allowing more critical video information to be preserved in the latent representation.
Variational Autoencoders.[18] introduced the VAE based on variational inference, establishing a novel generative network structure. Subsequent research [24, 34, 30] demonstrated that training and inference in VAE latent space could substantially reduce computational costs for diffusion models. [27] further proposed a two-stage image synthesis approach by decoupling perceptual compression from diffusion model. After that, numerous studies explored video VAEs with a focus on more efficient video compression including Open-Sora Plan [19], CogVideoX [39], and other models [3, 40, 46, 44, 6, 47]. Current video VAE architectures largely derive from earlier image VAE design [27, 10] and use a convolutional backbone. To address the challenges of inference on large videos, current video VAEs employ block-wise inference.
Latent Video Diffusion Models. In the early stages of Latent Video Diffusion Models (LVDMs) development, models like AnimateDiff [13], and MagicTime [41, 42] primarily utilized the U-Net backbone [28] for denoising, without temporal compression in VAEs. Following the paradigm introduced by Sora, recent open-source models such as Open-Sora [45], Open-Sora Plan [19, 20], and CogVideoX [39] have adopted DiT backbone with a spatiotemporally compressed VAE. Some methods including Open-Sora [45] and Movie Gen [25], employ a 2.5D design in either the DiT backbone or the VAE to reduce training costs. For LVDMs, the upper limit of video generation quality is largely determined by the VAE’s reconstruction quality. As overhead increase rapidly in scale, optimizing the VAE becomes crucial for handling large-scale data and enabling extensive pre-training.
@@ 3.1 Wavelet Transform
Preliminary. The Haar wavelet transform [23, 26], a fundamental form of wavelet transform, is widely used in signal processing [11, 12, 15]. It efficiently captures spatio-temporal information by decomposing signals through two complementary filters. The first is the Haar scaling filter $\mathbf{h}=\frac{1}{\sqrt{2}}[1,1]$, which acts as a low-pass filter that captures the average or approximation coefficients. The second is the Haar wavelet filter $\mathbf{g}=\frac{1}{\sqrt{2}}[1,-1]$, which functions as a high-pass filter that extracts the detail coefficients. These orthogonal filters are designed to be simple yet effective, with the scaling filter smoothing the signal and the wavelet filter detecting local changes or discontinuities.
@@
where $f_{i},f_{j},f_{k}\in\{\mathbf{h},\mathbf{g}\}$ represent the filters applied along each dimension, and * represents the convolution operation. The transform begins with $\mathbf{S}^{(0)}=\mathbf{V}$, and for subsequent layers, $\mathbf{S}^{(l)}=\mathbf{S}_{hhh}^{(l-1)}$, indicating that each layer operates on the low-frequency component from the previous layer. At each decomposition layer $l$, the transform produces eight sub-band components: $\mathcal{W}^{(l)}=\{\mathbf{S}_{hhh}^{(l)},\mathbf{S}_{hhg}^{(l)},\mathbf{S}_{hgh}^{(l)},\mathbf{S}_{ghh}^{(l)},\mathbf{S}_{hgg}^{(l)},\mathbf{S}_{ggh}^{(l)},\mathbf{S}_{ghg}^{(l)},\mathbf{S}_{ggg}^{(l)}\}$. Here, $\mathbf{S}_{hhh}^{(l)}$ represents the low-frequency component across all dimensions, while $\mathbf{S}_{ggg}^{(l)}$ captures high-frequency details. To implement different downsampling rates in the temporal and spatial dimensions, a combination of 2D and 3D wavelet transforms can be implemented. Specifically, to obtain a compression rate of 4×8×8 (temporal×height×width), we can employ a combination of two-layer 3D wavelet transform followed by one-layer 2D wavelet transform.
@@ 3.2 Architecture Design of WF-VAE
Through analyzing different sub-bands, we found that video energy is mainly concentrated in the low-frequency sub-band $\mathbf{S}^{(1)}_{hhh}$. Based on this observation, we establish an energy flow pathway outside the backbone, as show in Fig. 2, so that low-frequency information can smoothly flow from video to latent representation during encoding process, and then flow back to the video during the decoding process. This would inherently allow the model to attention more on low-frequency information, and apply higher compression rates to high-frequency information. With the additional path, we can reduce the computational cost brought by the dense 3D convolutions in the backbone.
Specifically, given a video $\mathbf{V}$, we apply multi-level wavelet transform to obtain pyramid features $\mathcal{W}^{(1)},\mathcal{W}^{(2)}$ and $\mathcal{W}^{(3)}$. We utilize $\mathcal{W}^{(1)}$ as the input for the encoder and the target output for the decoder. The convolutional backbone and multi-level wavelet transform layer downsample the feature maps simultaneously at every downsampling layer, enabling the concatenation of features from two branches in the backbone. We employ Inflow Block to transform the channel numbers of $\mathcal{W}^{(2)}$ and $\mathcal{W}^{(3)}$ to $C_{flow}$, which are then concatenated with feature maps from backbone. We compare $C_{flow}$ to the width of the energy flow pathway, as analyzed in Sec. 4.4. On the decoder side, we maintain a structure symmetrical to the encoder. We extract feature maps with $C_{flow}$ channels from the backbone and process them through Outflow Block to obtain ${\hat{\mathcal{W}}^{(2)},\hat{\mathcal{W}}^{(3)}}$. In order to allow information to flow from the lower layer to the $hhh$ sub-band of the next layer, we have:
@@
Overall, we create shortcuts for low-frequency information, ensuring its greater emphasis in latent representation.
@@ 3.3 Causal Cache
We replace regular 3D convolutions with causal 3D convolutions [40] in WF-VAE. The causal convolution applies $k_{t}-1$ temporal padding at the start with kernel size $k_{t}$. This padding strategy ensures the first frame remains independent from subsequent frames, thus enabling the processing of images and videos within a unified architecture. Furthermore, we leverage the causal convolution properties to achieve lossless inference. We first extract the initial frame from a video with $T$ frames. The remaining $T-1$ frames are then partitioned into temporal chunks, where $T_{chunk}$ represents the chunk size. Let $s_{t}$ denote the temporal convolutional stride, and $m=0,1,2,\cdots$ represents the chunk block index. To maintain the continuity of convolution sliding windows, each chunk caches its tail frames for the next chunk. The number of cached frames is given by:
@@ 3.4 Training Objective
The final loss function is formulated as:
@@ 4.1 Experimental Setup
Baseline Models. To assess the effectiveness of WF-VAE, we perform a comprehensive evaluation, comparing its performance and efficiency against several state-of-the-art VAE models. The models considered are: (1) OD-VAE [6], a 3D causal convolutional VAE used in Open-Sora Plan 1.2 [19]; (2) Open-Sora VAE [45]; (3) CV-VAE [44]; (4) CogVideoX VAE [39]; (5) Allegro VAE [46]; (6) SVD-VAE [4], which does not compress temporally and (7) SD-VAE [27], a widely used image VAE. Among these, CogVideoX VAE has a latent dimension of 16, while others use a latent dimension of 4. Notably, all VAEs except CV-VAE and SD-VAE, have been validated on LVDMs previously, making them highly representative for comparison.
@@
PSNR ()($\uparrow$)( ↑ ) SSIM()($\uparrow$)( ↑ ) LPIPS ()($\downarrow$)( ↓ ) FVD ()($\downarrow$)( ↓ ) PSNR ()($\uparrow$)( ↑ ) SSIM ()($\uparrow$)( ↑ ) LPIPS()($\downarrow$)( ↓ ) FVD ()($\downarrow$)( ↓ )
@@
Table 1: Quantitative metrics of reconstruction performance. Results demostrate that WF-VAE achieves state-of-the-art on reconstrcution performance comparing with other VAEs on WebVid-10M[2] and Panda70M[7] datasets. TCPR represents the token compression rate and Chn indicates the number of latent channels. The highest result is highlighted in bold, and the second highest result is underlined.
Figure 4: Computational performance of encoding and decoding. We evaluate the encoding and decoding time and memory consumption across 33 frames with 256×256, 512×512, and 768×768 resolutions (benchmark models without causal convolution are tested with 32 frames). WF-VAE surpasses other VAE models by a large margin on both inference speed and memory efficiency.
Dataset & Evaluation. We utilize the Kinetics-400 dataset [16] for both training and validation. For testing, we employ the Panda70M [7] and WebVid-10M [2] datasets. To comprehensively evaluate the model’s reconstruction performance, we select Peak Signal-to-Noise Ratio (PSNR) [14], Learned Perceptual Image Patch Similarity (LPIPS) [43], and Structural Similarity Index Measure (SSIM) [35] as primary evaluation metrics. Additionally, we use Fréchet Video Distance (FVD) [33] to assess visual quality and temporal coherence. To assess our model’s performance in generating results with the diffusion model, we utilize the UCF-101 [32] and SkyTimelapse [37] datasets for conditional and unconditional training 100,000 steps. Following [22, 31], we extract 16-frame clips of 2,048 videos to compute FVD16. Additionally, we evaluate the Inception Score (IS) [29] exclusively on the UCF-101 dataset, as suggested by [22]. We select Latte-L [22] as the denoiser. Since our primary focus is not on the final generative performance but rather on whether the latent spaces of various video VAEs facilitate effective diffusion model training, we chose not to use the higher-performing Latte-XL.

### v3

@@ 1 Introduction
The release of Sora [5], a video generation model developed by OpenAI, has pushed the boundaries of synthesizing photorealistic videos, drawing unprecedented attention to the field of video generation. Recent advancements in Latent Video Diffusion Models (LVDMs), such as Open-Sora Plan [19], Open-Sora [49], CogVideoX [40], EasyAnimate [39], Movie Gen [25], and ConsisID [44], have led to substantial improvements in video generation quality. These methods establish a compressed latent space [27] using a pre-trained video Variational Autoencoder (VAE), where the compression quality fundamentally determines the generative performance.
@@
Current video VAEs remain constrained by fully convolutional architectures inherited from the image era. They address video flickering and redundant information by incorporating spatio-temporal interaction and spatio-temporal compression layers. Several recent works, including OD-VAE [19, 6], CogVideoX [40], CV-VAE [48], and Allegro [50] adopt dense 3D structure to achieve high-quality video compression. While these methods demonstrate impressive reconstruction performance, they require prohibitively intensive computational resources. In contrast, alternative approaches such as Movie Gen [25] and Open-Sora [49] utilize 2+1D architecture, resulting in reduced computational requirements at the cost of lower reconstruction quality. Previous architectures have inadequately leveraged temporal redundancy in video data, necessitating the use of redundant spatio-temporal interaction layers to improve video compression quality.
@@
Many attempts employ block-wise inference strategies to trade computation time for memory, thus addressing computational bottlenecks in processing high-resolution, long-duration videos. EasyAnimate [39] introduced Slice VAE to encode and decode video frames in groups, but this approach leads to discontinuous output videos. Several methods, including Open-Sora [49], Open-Sora Plan [19], Allegro [50], implement tiling inference strategies but often produce spatio-temporal artifacts in overlapping regions. Although CogVideoX [40] employ caching to ensure convolution continuity, its reliance on group normalization [37] disrupts the independence of temporal feature, thus preventing lossless block-wise inference. These limitations highlight two core challenges faced by video VAEs: (1) excessive computational demands due to redundant architecture, and (2) compromised latent space integrity resulting from existing tiling inference strategies, which cause artifacts and flickering in reconstructed videos.
@@
Wavelet transform [23, 26] decomposes videos into multiple frequency-domain components. This decomposition enables prioritization strategies for encoding crucial video components. In this work, we propose Wavelet Flow VAE (WF-VAE), a novel autoencoder that utilizes multi-level wavelet transforms for extracting multi-scale pyramidal features and establishes a main energy flow pathway for these features to flow into latent representation. This pathway bypasses low-frequency video information to latent space, skipping the backbone network. Our WF-VAE enables a simplified backbone design with reduced 3D convolutions, significantly reducing computational costs. To address the potential latent space disruption, we propose Causal Cache mechanism. This approach leverages the properties of causal convolution. It maintains the continuity of the convolution sliding window through a caching strategy, which ensures numerical identity between block-wise inference and direct inference results. Experimental results show that WF-VAE achieves state-of-the-art reconstruction quality and computational efficiency performance. To summarize, major contributions of our work include:
@@
  • We introduce a lossless block-wise inference mechanism called Causal Cache, which maintains identical performance as direct inference across videos of any duration.
  • Extensive experimental evaluations of video reconstruction and generation demonstrate that WF-VAE achieves state-of-the-art performance.
@@ 2 Related Work
Figure 2: Overview of WF-VAE. Our architecture consists of a backbone and a main energy flow pathway. The pathway functions as a “highway” for the main flow of video energy, channeling this energy into the backbone through concatenations, allowing more critical video information to be preserved in the latent representation.
Variational Autoencoders.[18] introduced the VAE based on variational inference, establishing a novel generative network structure. Subsequent research [24, 35, 31] demonstrated that training and inference in VAE latent space could substantially reduce computational costs for diffusion models. [27] further proposed a two-stage image synthesis approach by decoupling perceptual compression from diffusion model. After that, numerous studies explored video VAEs with a focus on more efficient video compression, including Open-Sora Plan [19], CogVideoX [40], and other models [3, 41, 50, 48, 6, 51]. Current video VAE architectures primarily derive from earlier image VAE design [27, 9] and use a convolutional backbone. LiteVAE [29] employs a wavelet-based encoder inspired by latent-pixel similarity but overlooks energy perspectives, causing encoding-decoding energy asymmetry.
Latent Video Diffusion Models. In the early stages of Latent Video Diffusion Models (LVDMs) development, models like AnimateDiff [13], and MagicTime [45, 46] primarily utilized the U-Net backbone [28] for denoising, without temporal compression in VAEs. Following the paradigm introduced by Sora, recent open-source models such as Open-Sora [49], Open-Sora Plan [19, 20], and CogVideoX [40] have adopted DiT backbone with a spatiotemporally compressed VAE. Some methods, including Open-Sora [49] and Movie Gen [25], employ a 2+1D design in either the DiT backbone or the VAE to reduce training costs. For LVDMs, the upper limit of video generation quality is primarily determined by the VAE’s reconstruction quality. As overhead increases rapidly with scale, optimizing the VAE becomes essential for processing large-scale data and enabling extensive pre-training.
@@ 3.1 Wavelet Transform
Preliminary. The Haar wavelet transform [23, 26], a fundamental form of wavelet transform, is widely used in signal processing [11, 12, 15, 43, 42]. It efficiently captures spatio-temporal information by decomposing signals through two complementary filters. The first is the Haar scaling filter $\mathbf{h}=\frac{1}{\sqrt{2}}[1,1]$, which acts as a low-pass filter that captures the average or approximation coefficients. The second is the Haar wavelet filter $\mathbf{g}=\frac{1}{\sqrt{2}}[1,-1]$, which functions as a high-pass filter that extracts the detail coefficients. These orthogonal filters are designed to be simple yet effective, with the scaling filter smoothing the signal and the wavelet filter detecting local changes or discontinuities.
@@
where $f_{i},f_{j},f_{k}\in\{\mathbf{h},\mathbf{g}\}$ represent the filters applied along each dimension, and * represents the convolution operation. The transform begins with $\mathbf{S}^{(0)}=\mathbf{V}$, and for subsequent layers, $\mathbf{S}^{(l)}=\mathbf{S}_{hhh}^{(l-1)}$, indicating that each layer operates on the low-frequency component from the previous layer. At each decomposition layer $l$, the transform produces eight sub-band components: $\mathcal{W}^{(l)}=\{\mathbf{S}_{hhh}^{(l)},\mathbf{S}_{hhg}^{(l)},\mathbf{S}_{hgh}^{(l)},\mathbf{S}_{ghh}^{(l)},\mathbf{S}_{hgg}^{(l)},\mathbf{S}_{ggh}^{(l)},\mathbf{S}_{ghg}^{(l)},\mathbf{S}_{ggg}^{(l)}\}$. Here, $\mathbf{S}_{hhh}^{(l)}$ represents the low-frequency component across all dimensions, while $\mathbf{S}_{ggg}^{(l)}$ captures high-frequency details. To implement different downsampling rates in the temporal and spatial dimensions, a combination of 2D and 3D wavelet transforms can be implemented. Specifically, to obtain a compression rate of 4×8×8 (temporal×height×width), we can employ a combination of a two-layer 3D wavelet transform followed by a one-layer 2D wavelet transform.
@@ 3.2 Architecture Design of WF-VAE
Through analyzing different sub-bands, we find that video energy is mainly concentrated in the low-frequency sub-band $\mathbf{S}^{(1)}_{hhh}$. Based on this observation, we establish an energy flow pathway, as illustrated in Fig. 2, so that low-frequency information can smoothly flow from video to latent representation during the encoding process and then flow back to the video during the decoding process. This would inherently allow the model to pay more attention to low-frequency information and apply higher compression rates to high-frequency details. With the additional path, we can reduce the computational cost brought by the dense 3D convolutions in the backbone.
Specifically, given a video $\mathbf{V}$, we apply multi-level wavelet transform to obtain pyramid features $\mathcal{W}^{(1)},\mathcal{W}^{(2)}$ and $\mathcal{W}^{(3)}$. We utilize $\mathcal{W}^{(1)}$ as the input for the encoder and the target output for the decoder. The backbone and multi-level wavelet transform downsample the feature maps simultaneously at every downsampling layer, enabling the concatenation of features from two branches in the backbone. We employ Inflow Block to transform the channel numbers of $\mathcal{W}^{(2)}$ and $\mathcal{W}^{(3)}$ to $C_{flow}$, which are then concatenated with feature maps from backbone. We compare $C_{flow}$ to the width of the energy flow pathway, as analyzed in Sec. 4.3. On the decoder side, we maintain a structure symmetrical to the encoder. We split feature maps with $C_{flow}$ channels from the backbone and process them through Outflow Block to obtain ${\hat{\mathcal{W}}^{(2)},\hat{\mathcal{W}}^{(3)}}$. To allow information to flow from the lower layer to the $hhh$ sub-band of the next layer, we have:
@@
Overall, we strategically design shortcuts to prioritize low-frequency information, thereby enhancing its representation within the latent space.
@@ 3.3 Causal Cache
We replace regular 3D convolutions with causal 3D convolutions [41] in WF-VAE. The causal convolution applies $k_{t}-1$ temporal padding at the start with kernel size $k_{t}$. This padding strategy ensures that the first frame remains independent from subsequent frames, thus enabling the processing of images and videos within a unified architecture. Furthermore, we leverage the causal convolution properties to achieve lossless inference. We first extract the initial frame from a video with $T$ frames. The remaining $T-1$ frames are then partitioned into temporal chunks, where $T_{chunk}$ represents the chunk size. Let $s_{t}$ denote the temporal convolutional stride, and $m=0,1,2,\cdots$ represents the chunk block index. To maintain the continuity of convolution sliding windows, each chunk caches its tail frames for the next chunk. The number of cached frames is given by:
@@ 3.4 Training Objective
The final loss function is formulated as follows:
@@ 4.1 Experimental Setup
Baseline Models. To assess the effectiveness of WF-VAE, we perform a comprehensive evaluation, comparing its performance and efficiency against several state-of-the-art VAE models. The models considered are: (1) OD-VAE [6], a 3D causal convolutional VAE used in Open-Sora Plan 1.2 [19]; (2) Open-Sora VAE [49]; (3) CV-VAE [48]; (4) CogVideoX VAE [40]; (5) Allegro VAE [50]; (6) SVD-VAE [4], which does not compress temporally and (7) SD-VAE [27], a widely used image VAE. Among these, CogVideoX VAE adopts a latent dimension of 16, whereas others utilize a latent dimension of 4. Notably, most VAEs have been validated on LVDMs, making them highly representative for comparison.
@@
PSNR ()($\uparrow$)( ↑ ) SSIM()($\uparrow$)( ↑ ) LPIPS ()($\downarrow$)( ↓ ) rFVD ()($\downarrow$)( ↓ ) PSNR ()($\uparrow$)( ↑ ) SSIM ()($\uparrow$)( ↑ ) LPIPS()($\downarrow$)( ↓ ) rFVD ()($\downarrow$)( ↓ )
@@
Table 1: Quantitative metrics of reconstruction performance. Results demostrate that WF-VAE achieves state-of-the-art on reconstrcution performance comparing with other VAEs on WebVid-10M [2] and Panda70M [7] datasets. TCPR represents the token compression rate, and Chn indicates the number of latent channels. The highest result is highlighted in bold, and the second highest result is underlined.
Figure 4: Computational performance of encoding and decoding. We evaluate the encoding, decoding time, and memory consumption across 33 frames with 256×256, 512×512, and 768×768 resolutions (benchmark models without causal convolution are tested with 32 frames). WF-VAE surpasses other VAE models by a large margin in terms of both inference speed and memory efficiency.
Dataset & Evaluation. We utilize the Kinetics-400 dataset [16] for both training and validation. For testing, we employ the Panda70M [7] and WebVid-10M [2] datasets. To comprehensively evaluate the model’s reconstruction performance, we select Peak Signal-to-Noise Ratio (PSNR) [14], Learned Perceptual Image Patch Similarity (LPIPS) [47], and Structural Similarity Index Measure (SSIM) [36] as primary evaluation metrics. Additionally, we use reconstruction Fréchet Video Distance (rFVD) [34] to assess visual quality and temporal coherence. To assess our model’s performance in generating results with the diffusion model, we utilize the UCF-101 [33] and SkyTimelapse [38] datasets for conditional and unconditional training 100,000 steps. Following [22, 32], we extract 16-frame clips of 2,048 videos to compute FVD16. Additionally, we evaluate the Inception Score (IS) [30] exclusively on the UCF-101 dataset, as suggested by [22]. We select Latte-L [22] as the denoiser. Since we focus not on the generative performance but on whether the latent spaces of various video VAEs facilitate practical diffusion model training, we chose not to use the higher-performing Latte-XL.